# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

### v2

@@ Abstract
Recent advances in large reasoning models highlight Reinforcement Learning with Verifiable Rewards (RLVR) as a promising method for enhancing AI’s capabilities, particularly in solving complex logical tasks. However, it remains unclear whether the current practice of RLVR truly expands a model’s reasoning boundary or mainly amplifies high-reward outputs that the base model already knows for improved precision. This study presents an empirical investigation that provides fresh insights into the potential limits of the common practice of RLVR. We examine how, under current training conditions, RLVR can operate as a support-constrained optimization mechanism that may restrict the discovery of entirely original solutions, remaining constrained by the base model’s initial distribution. We also identify an entropy–reward trade-off: while the current RLVR recipe reliably enhances precision, it may progressively narrow exploration and potentially overlook correct yet underrepresented solutions. Extensive empirical experiments validate that while the current RLVR recipe consistently improves pass@1, the shrinkage of empirical support generally outweighs the expansion of empirical support under larger sampling budgets, failing to recover correct answers that were previously accessible to the base model. Interestingly, we also observe that while RLVR sometimes increases token-level entropy—resulting in greater uncertainty at each generation step—answer-level entropy declines, indicating that these seemingly more uncertain paths ultimately converge onto a smaller set of distinct answers. Taken together, these findings reveal potential limits of the current RLVR recipe in extending reasoning horizons. Breaking this invisible leash may require future algorithmic innovations such as explicit exploration mechanisms or hybrid strategies that seed probability mass into underrepresented solution regions.
@@ 1 Introduction
The rise of large reasoning models, such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024), marks a breakthrough in AI capabilities, particularly in solving complex logical tasks involving mathematics (Luo et al., 2025c; Zeng et al., 2025) and programming (Luo et al., 2025b; Liu & Zhang, 2025). The key ingredient behind this remarkable progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), where a pretrained base model—or one fine-tuned on long-form Chain-of-Thought (CoT) data—is optimized via reinforcement learning (RL) using simple, automatically computed rewards. Despite the empirical success, a fundamental question remains under active debate within the research community: does the current practice of RLVR expand a base model’s reasoning capabilities, or does it simply reinforce patterns the base model already knows, sometimes at the expense of exploring alternative correct solutions?
@@
Recent empirical studies have revealed a puzzling pattern that hints at this limitation. While models trained with the common RLVR recipe consistently outperform their base models when evaluated with a single attempt (pass@1), the base models often perform better when given multiple attempts (pass@k for large k). It is even reported that RLVR-trained models benefit from seemingly random or spurious reward signals, raising questions about whether the observed improvements genuinely reflect enhanced reasoning (Shao et al., 2025). While pass@k may have limitations as a comprehensive measure of reasoning boundaries—as it primarily captures solution retrieval rather than novel reasoning capacity (Wen et al., 2025)—we adopt it here as a practical proxy metric following established practice in prior work (Chen et al., 2021; Shao et al., 2024; Liu et al., 2024; Chen et al., 2025b). This metric provides a useful lens for examining how RLVR affects solution accessibility, though future work should explore more nuanced measures of reasoning capability expansion.
@@
Some researchers interpret this as evidence that the current RLVR recipe primarily performs conservative optimization within the base model’s existing capabilities (Yue et al., 2025a; Zhao et al., 2025; Shah et al., 2025; Ma et al., 2025; He et al., 2025). Others argue that this pattern only appears in specialized domains where base models were already well-trained, and that RLVR can substantially expand reasoning in other domains (Liu et al., 2025).
Figure 1: Empirical support dynamics of RLVR. Left: Conceptual illustration of empirical support under a threshold $\epsilon$. We define four regions based on whether a correct completion $y^{*}\in\mathcal{C}$ is assigned non-negligible probability mass by the base model $q$ and the RLVR model $\pi_{\theta}$: Empirical Support Preservation covers completions with $q(y^{*}|x)>\epsilon$ and $\pi_{\theta}(y^{*}|x)>\epsilon$; Empirical Support Shrinkage includes correct completions downweighted by RLVR below $\epsilon$; Empirical Support Expansion includes completions that RLVR newly upweights above $\epsilon$ despite negligible base model mass; and Out of Support refers to completions missed by both. Right: Pie charts showing the proportion of completions in each category across diverse reasoning tasks.
@@
Seeking a definitive answer to this debate remains an open challenge. In the extreme case, it seems unlikely that the current RLVR recipe can unlock advanced reasoning capabilities for any model out of the box, such as GPT-2 (Radford et al., 2019). We are curious if there may exist inherent limitations in the current RLVR practice. This paper provides a systematic empirical investigation into the fundamental capabilities and potential limitations of the current RLVR practice. We introduce the concept of empirical support—the set of correct solutions that a model can realistically discover under finite sampling. Using this framework, we show that:
@@
    The current RLVR recipe primarily preserves rather than expands the base model’s solution coverage. Across diverse reasoning benchmarks, RLVR consistently loses access to more correct solutions than it gains, even while improving single-sample accuracy.
@@
    The precision-diversity trade-off is fundamental, not domain-specific. This pattern appears across mathematics, logical reasoning, factual QA, and code generation—suggesting it reflects inherent properties of current RLVR methods rather than domain-specific quirks.
@@
These findings suggest that current RLVR methods may face an “invisible leash”—they remain fundamentally constrained by their initialization and cannot discover reasoning patterns that lie outside the base model’s effective reach. To break this invisible leash, RLVR may need augmenting with explicit exploration or hybrid strategies that seed probability mass into underrepresented regions of the solution space. We hope this work offers novel insights into the strengths and limitations of the current RLVR recipe, guiding future efforts in improving RLVR practice and building LLM systems that can unlock genuinely new reasoning capacity.
@@ 2.1 Formalizing Solution Accessibility
Let $\mathcal{X}$ denote the space of natural language prompts, and $\mathcal{Y}$ denote the space of token sequences (e.g., reasoning traces or completions). For a fixed prompt $x\in\mathcal{X}$, $q(y\mid x)$ is the output distribution of the base model, and $R(x,y)\in\{0,1\}$ is a verifiable reward function indicating whether $y$ is a correct solution. Various RLVR algorithms, including PPO (Schulman et al., 2017), RLOO (Kool et al., 2019), GRPO (Guo et al., 2025), DAPO (Yu et al., 2025), or REINFORCE++ (Hu, 2025), learn a new distribution $\pi_{\theta}(y\mid x)$ to optimize different variants of the following regularized objective:
$\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x),x\sim\mathcal{D}}\left[R(x,y)-\beta^{-1}\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\right],$
where $\mathcal{D}$ is the distribution of prompts. An optional log ratio corresponds to a regularized policy update that penalizes divergence from the base model $q$ controlled by a hyperparameter $\beta>0$.
@@ 3.1 Experimental Setup
Table 1: Aggregate support dynamics across diverse models and domains. Each completion is categorized by correctness and support status: Preservation indicates both base and RLVR find the solution; Shrinkage indicates the base model found it but RLVR did not; Expansion indicates only RLVR found it; and Out of Support denotes solutions found by neither. Higher SRR, NDR, and SDS reflect stronger preservation, genuine discovery, and balanced optimization, respectively. NSCR values closer to zero indicate more balanced support change. Kangheng-OVR-7B is included as a vision-language model (VLM). Full detailed statistics for each model are provided in Appx. A.
@@ 3.2 Results: Predominant Preservation with Limited Expansion
Perplexity analysis on support constraints. Tab. 2 reports perplexity scores under two complementary settings. When evaluated against external reasoning traces from DeepSeek-R1 and Claude Sonnet 4 with extended thinking—likely outside the base model’s support—RLVR shows markedly higher perplexity (e.g., AIME2024 rising from $8.76$ to $14.91$), confirming its inability to assign mass to fundamentally novel solution modes. Differences in language style and reasoning format across external references (e.g., Claude vs DeepSeek) also contribute to perplexity gaps, beyond purely structural support constraints. Breakdowns by correctness patterns highlight the precision–coverage trade-off: in shrinkage cases, ProRL’s perplexity rises when it fails to recover solutions still accessible to the base, reflecting entropy-driven concentration. Meanwhile, modest perplexity gaps in rare expansion cases indicate that these new completions were drawn from the base’s long-tail low-density regions—amplified but not truly beyond its support.
@@
Table 2: Perplexity of reasoning tokens from base and RLVR across math benchmarks, segmented by correctness patterns and reference types. Top: on problems correctly solved by the base model but not RLVR, perplexity is measured against the base model’s reasoning traces. Middle: on problems correctly solved by RLVR but not the base, perplexity is measured against RLVR’s traces. Bottom: on problems unsolved by both, perplexity is computed against external references (DeepSeek-R1 and Claude Sonnet 4), reflecting each model’s compatibility with broader solution modes.
@@
Shrinkage ✓ Base, ✗ ProRL Base Base 1.36 1.47 1.30
ProRL 1.60 1.84 1.50
Expansion ✗ Base, ✓ ProRL ProRL Base - - 1.52
ProRL - - 1.32
@@ Appendix A Detailed Statistics for Support Dynamics
Table 5: Support dynamics metrics and pass@k performance of ProRL-1.5B-v2, compared with its base model, DeepSeek-R1-Distill-Qwen-1.5B.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@8192)
AIME2024 93.3% 90.0% 0.964 0.000 0.000 -0.036 27 0 1 2
AIME2025 80.0% 73.3% 0.875 0.045 0.086 -0.080 21 1 3 5
AMC 100.0% 100.0% 1.000 0.000 0.000 0.000 40 0 0 0
Math 99.6% 99.6% 1.000 0.000 0.000 0.000 498 0 0 2
Minerva 71.7% 63.2% 0.856 0.029 0.056 -0.115 167 5 28 72
Olympiad 92.7% 88.7% 0.952 0.005 0.010 -0.043 596 3 30 46
Non-Math Reasoning Benchmarks (pass@2048)
SimpleQA 23.3% 22.2% 0.822 0.135 0.233 -0.044 83 13 18 319
LiveBench-R 100.0% 92.0% 0.920 0.000 0.000 -0.080 92 0 8 0
LiveBench-C 62.5% 60.2% 0.887 0.078 0.143 -0.035 71 6 9 42
LiveBench-L 26.0% 22.0% 0.692 0.182 0.288 -0.133 9 2 4 35
SciBench 94.1% 88.9% 0.937 0.008 0.016 -0.055 610 5 41 36
LiveCodeBench v5 46.4% 41.2% 0.820 0.075 0.138 -0.106 123 10 27 163
LiveCodeBench v6 43.5% 41.2% 0.895 0.056 0.105 -0.050 51 3 6 71
Aggregate Statistics
Math Benchmarks 0.956 0.007 0.013 -0.037 1349 9 62 127
Non-Math Benchmarks 0.902 0.036 0.070 -0.062 1039 39 113 666
Overall 0.932 0.020 0.039 -0.049 2388 48 175 793

### v3

@@ Abstract
Recent advances in large reasoning models highlight Reinforcement Learning with Verifiable Rewards (RLVR) as a promising method for enhancing AI’s capabilities, particularly in solving complex logical tasks. However, it remains unclear whether the current practice of RLVR truly expands a model’s reasoning boundary or mainly amplifies high-reward outputs that the base model already knows for improved precision. This study presents an empirical investigation that provides fresh insights into the potential limits of the common practice of RLVR. We examine how, under current training conditions, RLVR can operate as a support-constrained optimization mechanism that may restrict the discovery of entirely original solutions, remaining constrained by the base model’s initial distribution. We also identify an entropy–reward trade-off: while the current RLVR recipe reliably enhances precision, it may progressively narrow exploration and potentially overlook correct yet underrepresented solutions. Extensive empirical experiments validate that while the current RLVR recipe consistently improves pass@1, the shrinkage of empirical support generally outweighs the expansion of empirical support under larger sampling budgets, failing to recover correct answers that were previously accessible to the base model. Interestingly, we also observe that while RLVR sometimes increases token-level entropy, it results in greater uncertainty at each generation step and declining answer-level entropy. This indicates that these seemingly more uncertain paths ultimately converge onto a smaller set of distinct answers. Taken together, we reveal potential limits of the current RLVR recipe in extending reasoning horizons. Breaking this invisible leash may require future algorithmic innovations such as explicit exploration mechanisms or hybrid strategies that seed probability mass into underrepresented solution regions.
@@ 1 Introduction
The rise of large reasoning models, such as DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech and others, 2024), marks a breakthrough in AI capabilities, particularly in solving complex logical tasks involving mathematics (Luo et al., 2025c; Zeng et al., 2025) and programming (Luo et al., 2025b; Liu and Zhang, 2025). The key ingredient behind this remarkable progress is large-scale Reinforcement Learning with Verifiable Rewards (RLVR), where a pretrained base model or one fine-tuned on Chain-of-Thought data is optimized via RL using simple, automatically computed rewards. Prior work has also explored stronger notions of verifiability, such as models that can generate interactive proofs of their own correctness (Amit et al., 2025). Despite empirical success, a fundamental question remains under active debate within the research community: Does the current practice of RLVR expand base models’ reasoning capabilities, or simply reinforce patterns base models already know, sometimes at the expense of exploring alternative correct solutions?
@@
Recent studies have revealed a puzzling pattern that hints at this limitation. While models trained with the common RLVR recipe consistently outperform base models when evaluated with a single attempt (pass@1), base models often perform better given multiple attempts. It is even reported that RLVR benefits from seemingly random or spurious reward signals, raising questions about whether observed improvements genuinely reflect enhanced reasoning (Shao et al., 2025). While pass@k may have limitations as a comprehensive measure of reasoning boundaries, as it primarily captures solution retrieval rather than novel reasoning capacity (Wen et al., 2025), we adopt it here as a practical proxy metric following Chen et al. (2021); Shao and others (2024); Chen et al. (2025b). This metric provides a useful lens for examining how RLVR affects solution accessibility, though future work should explore more nuanced measures of reasoning capability expansion. Besides, prior work examines RLVR only through before/after snapshots, leaving unexplored how RLVR reshapes the model’s effective reasoning support throughout training.
@@
Some studies interpret pass@k as evidence that the current RLVR recipe primarily performs conservative optimization within the base model’s existing capabilities (Yue et al., 2025a; Zhao and others, 2025; Shah and others, 2025; Ma et al., 2025; He et al., 2025). Others argue that this pattern only appears in specialized domains where base models were already well-trained, and that RLVR can substantially expand reasoning in other domains (Liu et al., 2025).
@@
Seeking a definitive answer to this debate remains an open challenge. In the extreme case, it seems unlikely that the current RLVR recipe can unlock advanced reasoning capabilities for any model out of the box, such as GPT-2 (Radford et al., 2019). We are curious if there may exist inherent limitations in the current RLVR practice. This paper provides a systematic empirical investigation into the fundamental capabilities and potential limitations of the current RLVR practice. We introduce the concept of empirical support: the set of correct solutions that a model can realistically discover under finite sampling. Using this framework, we show that:
@@
    The current RLVR recipe primarily preserves rather than expands the base model’s solution coverage. Across diverse reasoning benchmarks, RLVR consistently loses access to more correct solutions than it gains, even while improving single-sample accuracy. We additionally analyze the temporal evolution of support dynamics across RLVR training steps, uncovering how RLVR progressively narrows the accessible solution space over time.
@@
    The precision-diversity trade-off is fundamental, not domain-specific. This pattern appears across mathematics, logical reasoning, factual QA, and code generation, suggesting it reflects inherent properties of current RLVR methods rather than domain-specific quirks.
@@
These findings show that current RLVR may face an “invisible leash”. They remain fundamentally constrained by their initialization and cannot discover reasoning patterns that lie outside the base model’s effective reach. To break this invisible leash, RLVR may need to be augmented with explicit exploration or hybrid strategies that seed probability mass into underrepresented regions of the solution space. We hope this work offers novel insights into the strengths and limitations of the current RLVR recipe, guiding future efforts in improving RLVR practice and building LLM systems that can unlock genuinely new reasoning capacity.
@@ 2.1 Formalizing Solution Accessibility
Let $\mathcal{X}$ denote the space of natural language prompts, and $\mathcal{Y}$ denote the space of token sequences (e.g., reasoning traces or completions). For a fixed prompt $x\in\mathcal{X}$, $q(y\mid x)$ is the output distribution of the base model, and $R(x,y)\in\{0,1\}$ is a verifiable reward function indicating whether $y$ is a correct solution. Various RLVR algorithms, including PPO (Schulman et al., 2017), RLOO (Kool et al., 2019), GRPO (Guo et al., 2025), DAPO (Yu et al., 2025), or REINFORCE++ (Hu, 2025), learn a new distribution $\pi_{\theta}(y\mid x)$ to optimize different variants of the following regularized objective: $\max_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x),x\sim\mathcal{D}}\left[R(x,y)-\beta^{-1}\log\frac{\pi_{\theta}(y\mid x)}{q(y\mid x)}\right],$ where $\mathcal{D}$ is the distribution of prompts. An optional log ratio corresponds to a regularized policy update that penalizes divergence from the base model $q$ controlled by a hyperparameter $\beta>0$.
@@ 3.1 Experimental Setup
Table 1: Aggregate support dynamics across diverse models and domains. Each completion is categorized by correctness and support status: Preservation indicates both base and RLVR find the solution; Shrinkage indicates only the base model found it; Expansion indicates only RLVR found it; and Out of Support denotes solutions found by neither. Higher SRR, NDR, and SDS reflect stronger preservation, genuine discovery, and balanced optimization, respectively. NSCR values closer to zero indicate more balanced support change. Kangheng-OVR-7B is included as a vision-language model (VLM). Full detailed statistics for each model are provided in Appx. A.
@@ 3.2 Results: Predominant Preservation with Limited Expansion
Perplexity analysis on support constraints. Tab. 3 reports perplexity where base and ProRL models are evaluated against external reasoning traces from DeepSeek-R1 and Claude Sonnet 4. ProRL consistently shows higher perplexity. For instance, on AIME2024, perplexity against Claude Sonnet 4 rises from 8.76 (Base) to 14.91 (ProRL), which indicates RLVR reduces the model’s ability to assign probability to diverse external reasoning styles. While format differences contribute, the dominant effect is structural: RLVR concentrates probability around narrower solution trajectories. Correctness patterns further reveal the precision-coverage trade-off. In shrinkage cases, ProRL shows higher perplexity even when the base model succeeds, confirming RLVR collapses mass away from viable solution pathways. Conversely, modest perplexity improvements in rare expansion cases suggest ProRL’s new successful trajectories originate from the base model’s low-density tails rather than genuinely novel reasoning structures.
Table 2: Comparison of support dynamics between SFT and DAPO training on Qwen2.5-Math-7B. SFT moderately expands support, while DAPO sharpens support but often reduces stability and increases shrinkage (lower NSCR).
Dataset Training pass@k Support Dynamics Metrics Support Counts
Base Target SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@256)
AIME2024 SFT 63.33% 63.33% 0.789 0.211 0.332 0.000 15 4 4 7
DAPO 63.33% 66.67% 0.895 0.150 0.257 0.045 17 3 2 8
AIME2025 SFT 50.00% 63.33% 0.933 0.263 0.411 0.200 14 5 1 10
DAPO 50.00% 56.67% 1.000 0.118 0.211 0.118 15 2 0 13
AMC23 SFT 100.00% 100.00% 1.000 0.000 0.000 0.000 40 0 0 0
DAPO 100.00% 92.50% 0.925 0.000 0.000 -0.075 37 0 3 0
MATH500 SFT 96.40% 99.00% 0.996 0.030 0.059 0.026 480 15 2 3
DAPO 96.40% 95.20% 0.975 0.013 0.025 -0.012 470 6 12 12
Minerva SFT 63.24% 66.91% 0.895 0.154 0.263 0.050 154 28 18 72
DAPO 63.24% 55.15% 0.802 0.080 0.145 -0.120 138 12 34 88
Olympiad SFT 78.22% 81.78% 0.953 0.089 0.162 0.042 503 49 25 98
DAPO 78.22% 72.89% 0.884 0.051 0.096 -0.065 467 25 61 122
Comparison between SFT and RLVR. We fix the base model (Qwen2.5-Math-7B), dataset (DeepMath-103K), sampling protocol, and optimization hyperparameters, varying only the training objective (SFT vs. DAPO). As shown in Tab. 2 and 12, SFT consistently produces moderate support expansion with positive NSCR values across benchmarks, whereas DAPO produces sharply concentrated distributions with mixed or negative support change. While DAPO improves precision at low sampling budgets, its support dynamics reflect a high-SRR but low-NDR regime: preservation of known correct solutions but limited discovery of new ones. On several benchmarks such as MATH500, Minerva, and Olympiad, DAPO reduces accessible correct solutions, causing net shrinkage despite identical training conditions. These results confirm that support-constrained behavior is not an artifact of scaling, data mixture, or procedure, but emerges from the objective itself. By contrast, SFT expands the reachable solution set without collapsing existing modes, underscoring that shrinkage is intrinsic to RLVR-style multiplicative updates rather than SFt.
Support dynamics during RLVR training. Tab, 13 summarizes how support-dynamics metrics evolve across RLVR training steps. Across benchmarks, SRR remains consistently high, showing that RLVR reliably preserves previously correct reasoning paths, while NSCR gradually decreases, indicating a steady contraction of the model’s solution support. On harder datasets, mild fluctuations in NDR and SDS reflect transient exploration that is subsequently pruned. Overall, these checkpoint-level statistics reveal that RLVR reshapes rather than monotonically improves the reasoning distribution. It reinforces a narrow set of stable trajectories over time, which explains both the early-stage gains and the late-stage degradation observed on more diverse math tasks.
@@
Table 3: Perplexity of reasoning tokens from base and RLVR across math benchmarks, segmented by correctness patterns and reference types. For different problem categories (e.g., shrinkage and expansion), perplexity is computed against external references (DeepSeek-R1 and Claude Sonnet 4), reflecting each model’s compatibility with diverse and broad solution modes.
@@
Shrinkage ✓ Base, ✗ ProRL DeepSeek-R1 Base 1.24 1.39 1.17
ProRL 1.39 1.70 1.25
Claude Sonnet 4 Base 1.70 1.54 1.51
ProRL 2.12 1.98 1.83
Expansion ✗ Base, ✓ ProRL DeepSeek-R1 Base - - 1.41
ProRL - - 1.28
Claude Sonnet 4 Base - - 1.65
ProRL - - 1.38
@@
3.3 When Empirical Support Expansion Occurs
Although rare, empirical support expansion follows clear patterns that align with the base model’s latent capabilities. Across Reasoning Gym tasks, we identify two primary mechanisms that explain why expansion arises in a small set of cases such as dice, arc_1d, boxnet, and graph_color_vertex20.
(1) RLVR Recomposes Subskills the Base Model Already Possesses
These expansion tasks exhibit modular or compositional structure, such as local moves in graph coloring, element-wise updates in arc_1d, or JSON-style key-value fragments in boxnet. The base model assigns non-negligible mass to many of these fragments individually, but not to their correct global combination. RLVR magnifies these weakly represented components and helps the model assemble them coherently. This matches our empirical patterns: all observed expansions occur in tasks where the base model’s pass@1 is low, but its pass@k curve rises steadily (Fig. 2), indicating that all fragments needed for correct reasoning already lie in the base model’s long tail. RLVR amplifies these long-tail valid completions, elevating a few above the empirical support threshold $\epsilon$.
(2) RLVR Corrects Prompt-Format Misalignment and Extracts Latent Ability
A second driver is the sensitivity to prompt format. In several expanding tasks, especially dice and boxnet, the base model demonstrates partial competence but fails to follow the exact response format required by the reward function. RLVR reshapes the distribution to follow instructions more effectively, unlocking capabilities that were previously present but suppressed. Consistent with this interpretation, the perplexity gaps in Table 3 remain modest in expansion cases, showing that the “new” completions are stylistic or formatting variants of reasoning patterns already accessible to the base model, not fundamentally novel solutions beyond its support.
Why Expansion Is Small and Bounded
Even when these two favorable mechanisms align, expansion remains sharply limited. Across all benchmarks, NDR never exceeds $0.04$ (Tab. 1), NSCR remains negative in all models. Thus, RLVR’s gains arise from amplifying low-probability but existing solution fragments or format-correct variants, but never from discovering solutions truly absent from the base distribution.
@@ 5 Conclusion
Acknowledgment
This work was supported in part by RS-2024-00457882, National AI Research Lab Project.
@@ Appendix A Detailed Statistics for Support Dynamics
Table 11: Support dynamics metrics and pass@256 performance of OLMo-2-0425-1B-DPO vs RLVR.
Dataset pass@k Performance Support Dynamics Metrics Support Counts
Base RLVR SRR NDR SDS NSCR P E S O
Math Reasoning Benchmarks (pass@256)
AIME2024 20.00% 16.67% 0.833 0.000 0.000 -0.167 5 0 1 24
AIME2025 23.33% 20.00% 0.571 0.333 0.421 -0.111 4 2 3 21
AMC23 85.00% 77.50% 0.912 0.000 0.000 -0.088 31 0 3 6
MATH500 77.80% 76.20% 0.920 0.060 0.113 -0.019 358 23 31 88
Minerva-Math 33.82% 32.35% 0.837 0.125 0.218 -0.039 77 11 15 169
Olympiad 49.93% 49.33% 0.849 0.141 0.242 -0.010 286 47 51 291
Aggregate Statistics
Math Benchmarks 0.887 0.110 0.166 -0.072 761 83 104 599
Table 12: Pass@k comparison of Qwen2.5-Math-7B (SFT vs DAPO) across five math benchmarks. Higher values indicate better performance.
AIME2024 @1 @2 @4 @8 @16 @32 @64 @128 @256
SFT 20.00 30.00 30.00 33.33 43.33 50.00 63.33 63.33 63.33
DAPO 20.00 26.67 26.67 40.00 46.67 56.67 60.00 63.33 66.67
AIME2025 @1 @2 @4 @8 @16 @32 @64 @128 @256
SFT 16.67 16.67 26.67 30.00 33.33 40.00 43.33 50.00 63.33
DAPO 16.67 20.00 30.00 43.33 46.67 50.00 50.00 50.00 56.67
AMC23 @1 @2 @4 @8 @16 @32 @64 @128 @256
SFT 60.00 72.50 77.50 85.00 90.00 95.00 97.50 97.50 100.00
DAPO 62.50 72.50 85.00 87.50 87.50 92.50 92.50 92.50 92.50
MATH500 @1 @2 @4 @8 @16 @32 @64 @128 @256
SFT 78.20 85.20 90.00 92.80 95.00 96.20 98.00 98.80 99.00
DAPO 85.00 89.80 91.40 93.00 93.20 93.80 94.40 94.80 95.20
Minerva-Math @1 @2 @4 @8 @16 @32 @64 @128 @256
SFT 36.03 39.71 43.75 47.79 51.47 54.04 63.60 65.44 66.91
DAPO 40.81 45.96 48.16 51.10 51.10 52.57 53.68 54.41 55.15
Table 13: Effect of RLVR (DAPO) training on DeepSeek 1.5B across math benchmarks. Pass@256 values are percentages. Step = 0 corresponds to the non-RLVR baseline.
Dataset Step Pass@256 Support Dynamics Metrics Support Counts
RLVR SRR NDR SDS NSCR P E S O
AIME24 0 83.33 - - - - - - - -
30 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
60 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
90 83.33 1.000 0.000 0.000 0.000 25 0 0 5
120 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
150 80.00 0.960 0.000 0.000 -0.040 24 0 1 5
180 83.33 1.000 0.000 0.000 0.000 25 0 0 5
210 76.67 0.920 0.000 0.000 -0.080 23 0 2 5
240 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
270 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
300 73.33 0.880 0.000 0.000 -0.120 22 0 3 5
AIME25 0 70.00 - - - - - - - -
30 70.00 0.905 0.095 0.172 0.000 19 2 2 7
60 73.33 0.905 0.136 0.237 0.042 19 3 2 6
90 70.00 0.905 0.095 0.172 0.000 19 2 2 7
120 73.33 0.905 0.136 0.237 0.042 19 3 2 6
150 73.33 0.905 0.136 0.237 0.042 19 3 2 6
180 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
210 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
240 66.67 0.905 0.050 0.095 -0.045 19 1 2 8
270 63.33 0.905 0.000 0.000 -0.095 19 0 2 9
300 66.67 0.864 0.095 0.172 -0.042 18 2 3 7
AMC23 0 97.50 - - - - - - - -
30 97.50 1.000 0.000 0.000 0.000 39 0 0 1
60 97.50 1.000 0.000 0.000 0.000 39 0 0 1
90 97.50 1.000 0.000 0.000 0.000 39 0 0 1
120 97.50 1.000 0.000 0.000 0.000 39 0 0 1
150 97.50 1.000 0.000 0.000 0.000 39 0 0 1
180 97.50 1.000 0.000 0.000 0.000 39 0 0 1
210 97.50 1.000 0.000 0.000 0.000 39 0 0 1
240 97.50 1.000 0.000 0.000 0.000 39 0 0 1
270 97.50 1.000 0.000 0.000 0.000 39 0 0 1
300 97.50 1.000 0.000 0.000 0.000 39 0 0 1
MATH500 0 99.20 - - - - - - - -
30 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
60 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
90 99.20 0.998 0.002 0.004 0.000 495 1 1 3
120 99.00 0.996 0.002 0.004 -0.002 494 1 2 3
150 98.80 0.996 0.000 0.000 -0.004 494 0 2 4
180 99.00 0.998 0.000 0.000 -0.002 495 0 1 4
210 98.80 0.994 0.002 0.004 -0.004 493 1 3 3
240 98.60 0.992 0.002 0.004 -0.006 492 1 4 3
270 98.40 0.992 0.000 0.000 -0.008 492 0 4 4
300 97.80 0.984 0.002 0.004 -0.014 488 1 8 3
Minerva 0 62.50 - - - - - - - -
30 62.13 0.935 0.059 0.111 -0.006 159 10 11 92
60 59.93 0.929 0.031 0.059 -0.040 158 5 12 97
90 61.76 0.924 0.065 0.122 -0.011 157 11 13 91
120 62.87 0.953 0.053 0.100 0.006 162 9 8 93
150 62.50 0.947 0.053 0.100 0.000 161 9 9 93
180 59.56 0.906 0.049 0.094 -0.045 154 8 16 94
210 60.29 0.924 0.043 0.082 -0.034 157 7 13 95
240 61.40 0.941 0.042 0.080 -0.017 160 7 10 95
270 59.19 0.906 0.043 0.083 -0.051 154 7 16 95
300 58.46 0.912 0.025 0.049 -0.063 155 4 15 98
Olympiad 0 88.00 - - - - - - - -
30 86.67 0.963 0.022 0.043 -0.015 572 13 22 68
60 86.81 0.960 0.027 0.053 -0.013 570 16 24 65
90 86.07 0.961 0.017 0.034 -0.022 571 10 23 71
120 86.52 0.961 0.022 0.044 -0.016 571 13 23 68
150 85.48 0.955 0.017 0.034 -0.028 567 10 27 71
180 84.00 0.944 0.011 0.021 -0.045 561 6 33 75
210 82.81 0.926 0.016 0.032 -0.058 550 9 44 72
240 82.81 0.931 0.011 0.021 -0.058 553 6 41 75
270 81.33 0.912 0.013 0.025 -0.075 542 7 52 74
300 79.70 0.897 0.009 0.018 -0.093 533 5 61 76